the preimeter of arectangle is 122inches. The length exceeds the width by 57 inches. find the length and the with?.
the length is?
the width is?

Answer by mananth(16946)

let the widtth be x
the length will be x+57
perimeter = 2*(L+W)
2*(x+x+57)=122
2*(2x+57)=122
4x+114=122
add-114 to both sides
4x+114-114=122-114
4x=8
divide by 4
4x/4 = 8/4
x=2 in- the width
length = x+57= 2+57=59in
..